Skip to content

Commit 12d846d

Browse files
authored
add entries for MD16906 and MD17423 (#221)
1 parent 01390dd commit 12d846d

File tree

1 file changed

+5
-0
lines changed

1 file changed

+5
-0
lines changed

docs/logbook/LHC/2025_lhc.md

Lines changed: 5 additions & 0 deletions
Original file line numberDiff line numberDiff line change
@@ -49,6 +49,8 @@ The tables below can be sorted by clicking next to the column headers.
4949
| 2025-11-02 23:00 | 2025-11-02 23:59 | 0.15HN | END-of-FILL | Luminosity measurement while trimming waist | [Shift Start](logbook://2025-11-02,1081,4435672) / [Summary](logbook://2025-11-02,1081,4441388) ([OP][op_lumitest]{.cern_login target=_blank}) |
5050
| 2025-11-04 22:30 | 2025-11-05 03:40 | 0.5WN | MD17443 | ADT-ACD measurements in OP config (indiv,MO,Q') | [Shift Start](logbook://2025-11-04,1081,4439816) / [Summary](logbook://2025-11-05,1081,4440023) ([OP][op_opconfig]{.cern_login target=_blank}) |
5151
| 2025-11-06 03:00 | 2025-11-06 09:30 | 0.8WN | MD17383 | New IR7 optics, squeeze validation | [Shift Start](logbook://2025-11-06,1081,4441445) / [Summary](logbook://2025-11-06,1081,4441726) ([OP][op_ir7squeeze]{.cern_login target=_blank}) |
52+
| 2025-11-07 00:01 | 2025-11-07 11:30 | 1WN 0.25W | MD16906 | Extra squeeze steps | [Shift Start](logbook://2025-11-07,1081,4442872) / [Summary](logbook://2025-11-07,1081,4443242) ([OP][op_md16906]{.cern_login target=_blank}) |
53+
| 2025-11-07 11:30 | 2025-11-07 13:30 | 0.25W | MD17423 | Identifying beam energy discrepancy | [Shift Start](logbook://2025-11-07,1081,4442872) / [Summary](logbook://2025-11-07,1081,4443242) |
5254
<!-- Logbook Links: [LINK_NAME](logbook://date, logbook_id, event_id) -->
5355

5456
<!-- OP logbook links below -->
@@ -78,6 +80,7 @@ The tables below can be sorted by clicking next to the column headers.
7880
[op_lumitest]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4436213
7981
[op_opconfig]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4440024
8082
[op_ir7squeeze]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4441728
83+
[op_md16906]: https://be-op-logbook.web.cern.ch/elogbook-server/GET/showEventInLogbook/4443198
8184
<!-- Tooltips -->
8285

8386
*[Injection Optics Global Corrs]: Test of tools. Rephased BPMs. Confirmed local coupling corrections, measured virgin machine at injection optics and did global corrections. Reached similar performance as previous years. First attempts of total phase corrections, used the ones from 2024. First use of new Segment-by-Segment.
@@ -104,5 +107,7 @@ The tables below can be sorted by clicking next to the column headers.
104107
*[Pb Ion Optics Part 3]: Attempted to validate Pb optics with enf of squeeze corrections from last time, as filler during validation of loss maps during vaccuum issue. Quick measurement at injection to check BPMs, but then OP were unable to get cycle to ramp due to inconsistency in ADT settings. Gave up after several hours as LHC switching back to lumi production. Includes summary of previous ion fill results.
105108
*[LRBB Beta-beat compensation with flat-optics]: Tried for the first time to take measurements of the LRBB-optics in both beam-1 and beam-2, using 2 separate fills. No issues encountered when switching the role of weak/strong beams between beam1/2. In Beam1 there was no visible effect on the lifetime. In beam2 however the lifetime appeared to degrade slightly as the MO were removed.
106109
*[New IR7 optics, squeeze validation]: Checked the new IR7 optics along the squeeze down to 60cm/18cm, including new corrections scaled from nominal cycle. Beta-beating and normalised dispersion good at all points in the squeeze. Vertical dispersion bump observed in IR7, getting stronger as the squeeze progresses (up to Dy=1m). No issue for the second fill of MD with lossmaps but should be corrected for operational use. The sdisp knob is to be included in model. Kmodulation at 18cm shows similar results to nominal ion cycle.
110+
*[Extra squeeze steps]: First part, tested a new squeeze down to 40cm/12cm. Stopped at 50cm/15cm where the optics were very good (similar to 60cm/18cm). At 40cm/12cm a peak beat-beating of 20% was observed (acceptable, should be corrected) and small waist shift. Unforeseen negative effects of Stephane's correction knobs brought peak beta-beating to 100% in the arcs and betastar to ~7-9cm. Did not proceed with fill of nominal bunches.
111+
*[Identifying beam energy discrepancy]: Benchmark data at 40cm/12cm with various DP_TRIM_PERMIL values to be analyzed offline.
107112

108113
--8<-- "docs/logbook/footer_shifts"

0 commit comments

Comments
 (0)